What is the molecular formula of a compound that is composed of 0.186 g C, 0.0390 g H

and 0.108 g N that has a molar mass of 129 g/mol?
Answer #1

Given:

weights in grams:

Carbon - 0.186 g

Hydrogen - 0.0390 g

Nitrogen - 0.108 g

Molar mass - 129 g/mol

i) Convert grams to moles

Carbon = g carbon/atomic weight = 0.186/12.0 = 0.0155

hydrogen = g hydrogen/ atomic weight = 0.0390/ 1 = 0.0390

Nitrogen = g nitrogen/ atomic weight = 0.108/14 = 0.0077

ii) Divide the number of moles of each element by the lowest number of mole i.e. 0.0077

Carbon = 0.0155/0.0077 = 2.01

Hydrogen = 0.0390/0.0077 = 5.06

Nitrogen = 0.0077/0.0077 = 1

iii) Round off the decimals to integers.

Carbon = 2

Hydrogen = 5

Nitrogen = 1

iv) The empirical formula : C2H5N1 , molar mass= 43

v) Molecular formula : (C2H5N1)n

n= (molar mass of the given compound)/(molar mass of empirical formula) = 129/43 = 3

Therefore, (C2H5N1)n = (C2H5N1)3 = C6H15N3

Common names = Acetaldehyde Ammonia Trimer, Aminoethylpiperazine.